### v1

# Cosmic Ray Superdiffusion and Mirror Diffusion in Partially Ionized and Turbulent Medium

Abstract
Understanding cosmic ray (CR) diffusion in a partially ionized medium is both crucial and challenging. In this study, we investigate CR superdiffusion in turbulent, partially ionized media using high-resolution 3D two-fluid simulations that treat ions and neutrals separately. We examine the influence of neutral-ion decoupling and the associated damping of turbulence on CR propagation in both transonic and supersonic conditions. Our simulations demonstrate that neutral-ion decoupling significantly damps velocity, density, and magnetic field fluctuations at small scales, producing spectral slopes steeper than those of Kolmogorov and Burgers scaling. We also identify an intermediate coupling regime in which neutrals remain partially coupled with ions, leading to a steepening of kinetic energy spectra in both fluids. Shocks can enhance the neutral-ion coupling, thereby reducing the differences between neutral and ion density structures. Moreover, the damping of magnetic field fluctuations substantially decreases pitch-angle scattering, which increases CR parallel mean free paths. As a result, CR perpendicular transport transitions between two distinct superdiffusive regimes: a scattering-dominated regime with perpendicular displacement proportional to $t^{3/4}$, and a scattering-free regime dominated by magnetic field line wandering, with perpendicular displacement scaling as $t^{3/2}$. When the pitch angle is large, the effects of magnetic mirroring—naturally arising in magnetohydrodynamic turbulence—become significant, enhancing the confinement (but not fully trapping) of CRs and inducing oscillations in their perpendicular displacement. These results highlight the necessity of incorporating two-fluid effects for accurately modeling CR transport in partially ionized environments such as molecular clouds and dense interstellar clumps.
Interstellar medium (847) — Cosmic rays (329) — Plasma astrophysics (1261) — Magnetohydrodynamics (1964)
software: Python3 (Van Rossum & Drake, 2009)
1 Introduction
[…]
A key determinant of CR transport is the nature of turbulent magnetic fields. Fluctuating fields introduce pitch-angle scattering and govern the effective diffusion of CRs along field lines (Jokipii, 1966; Jokipii & Parker, 1969; Qin et al., 2002; Lazarian & Xu, 2021). A growing body of work has revealed that CR transport can deviate significantly from the classical diffusion paradigm in the presence of magnetohydrodynamic (MHD) turbulence (Comisso & Sironi, 2018, 2019; Beattie et al., 2022; Lazarian et al., 2023; Zhang & Xu, 2024; Marcowith, 2025). In particular, turbulent magnetic field wandering (Lazarian & Vishniac (1999)) induces superdiffusive transport perpendicular to mean magnetic field (e.g., Yan & Lazarian, 2008; Xu & Yan, 2013; Lazarian & Yan, 2014; Hu et al., 2022), which in combination with the effects of CR bouncing in magnetic mirrors (Cesarsky & Kulsrud, 1973) results in parallel propagation that strongly depends on the CR pitch angle, i.e., in the process termed ”mirror diffusion” (Lazarian & Xu, 2021; Xu & Li, 2023). However, many studies assume a fully ionized plasma or utilize single-fluid approximations. In dense clumps, on the contrary, the gas is cold ($\sim$10–100 K) and only weakly ionized, resulting in additional physical processes such as ion-neutral collisions, charge exchange and turbulence damping (e.g., Draine, 1986; Balsara, 1996; Tilley & Balsara, 2011; Xu et al., 2015). These effects can alter both the local turbulence spectrum and CR scattering rates (Tilley & Balsara, 2010; McKee et al., 2010; Burkhart et al., 2015; Hu et al., 2024). Consequently, a two-fluid (ion + neutral) description is critical to fully capture the microphysics that governs CR propagation in these environments. By incorporating two-fluid effects, one can better connect observations—such as locally enhanced gamma-ray emission in molecular clouds—to the underlying plasma and magnetic field conditions. Consequently, a two-fluid (ion + neutral) description is critical to fully capture the microphysics that governs CR propagation in these environments. Incorporating two-fluid effects can better connect observations, such as locally enhanced gamma-ray emission in molecular clouds, to the underlying plasma and magnetic field conditions.
Despite the recognized importance of two-fluid processes, direct numerical simulations of this regime remain challenging because the need to resolve large Alfvén speeds in the neutral-ion decoupling layers quickly inflates computational costs. Earlier work has shown that neutral-ion decoupling can significantly modify the turbulent energy spectra of ions and neutrals (Tilley & Balsara, 2010; McKee et al., 2010; Burkhart et al., 2015). In particular, Hu et al. (2024) found that large density fluctuations in ions and neutrals can drive substantial spatial variations in the ionization fraction and local neutral-ion decoupling scale, leading to a globally intermediately coupled status. In this intermediately coupled regime, turbulence in both fluids can be damped, resulting in steeper kinetic energy spectra than the familiar Kolmogorov profile, which profoundly influences CR propagation.
In this work, we extend existing theories of CR transport and magnetic turbulence by incorporating essential two-fluid physics of neutral-ion decoupling and turbulence damping. We perform two-fluid MHD simulations using the AthenaK code (Stone et al., 2024), varying the ionization fraction and coupling strength to explore different regimes of ion-neutral interaction. By examining the resulting CR scattering properties and superdiffusive behavior, we aim to clarify how partial ionization modifies turbulence and, consequently, CR transport in dense interstellar environments.
This paper is organized as follows. § 2 briefly review the theoretical consideration of CR superdiffusion in a magnetized and partially ionized medium. § 3 describes the 3D two-fluid simulations and numerical approach of modeling CR propagation used in this study. In § 4, we analyze the statistics of turbulence and CR diffusion in the partially ionized medium across different neutral-ion coupling statuses. We conclude with a summary of our findings in § 5.
2 Theoretical consideration
2.1 Decoupling of neutrals and ions
The coupling between ions and neutrals is commonly characterized by two collisional frequencies. The neutral–ion collision frequency is defined as $\nu_{ni}=\gamma_{\rm d}\rho_{i}=\gamma_{\rm d}\xi_{i}(\rho_{i}+\rho_{n})$ while the ion–neutral collision frequency is given by $\nu_{in}=\gamma_{\rm d}\rho_{n}$, respectively (Shu, 1992). Decoupling of neutrals from ions begins when the rate at which turbulent energy cascades becomes comparable to $\nu_{ni}$. Because $\nu_{in}$ greatly exceeds $\nu_{ni}$ in weakly ionized media, ions separate from neutrals on much smaller scales than those marking the decoupling of neutrals; hence, our focus here is primarily on the neutral–ion decoupling process.
Early linear studies proposed that neutrals decouple from Alfvénic oscillations at a characteristic wavenumber $k_{{\rm dec},\parallel}$, determined by setting the Alfvén wave frequency equal to the neutral–ion collision frequency (Kulsrud & Pearce, 1969; Shu, 1992):
$k_{{\rm dec},\parallel}v_{A}=\nu_{ni}.$ (1)
Here, the subscript $\parallel$” designates the component along the magnetic field, and $v_{A}$ is the Alfvén speed.
Although MHD turbulence was once approximated as a superposition of linear waves (Giacalone & Jokipii, 1999), it is now recognized as an inherently non-linear process. In particular, within the strong turbulence regime under a critical balance condition, any wavelike behavior of Alfv’en is short-lived, persisting for less than a wave period (Goldreich & Sridhar, 1995; Lazarian & Vishniac, 1999). With the dynamically coupled turbulent mixing motion in the perpendicular direction and the wave-like motion in the parallel direction, the scale-dependent anisotropy of MHD turbulence develops. For Alfvénic turbulence  [^1] For a more in-depth discussion on neutral-ion decoupling scales of the three MHD modes (Alfvén, fast, and slow), see Xu et al. (2015, 2016)., which typically dominates the turbulent energy budget (Cho et al., 2002; Hu et al., 2021), this anisotropy implies that the decoupling length is itself anisotropic—the parallel component can exceed the perpendicular one, so that the latter is smaller than earlier estimates (Xu et al., 2015, 2016; Hu et al., 2024). The perpendicular decoupling scale in strong MHD turbulence regimes is expressed as (Xu et al., 2015):
$k_{{\rm dec},\bot}$ =$=\begin{cases}\nu_{\rm ni}^{3/2}L_{\rm inj}^{1/2}v_{\rm inj}^{-3/2},~{}~{}~{}M_{A}>1\\ \nu_{\rm ni}^{3/2}L_{\rm inj}^{1/2}v_{\rm inj}^{-3/2}M_{A}^{-1/2},~{}~{}~{}M_{A}<1\\ \end{cases},$ (2)
where $L_{\rm inj}$ and $v_{\rm inj}$ are the injection scale and injection velocity of turbulence, respectively. $M_{A}=v_{\rm inj}/v_{A}$ is the Alfvén Mach number.
Hu et al. (2024) further revealed that, in a turbulent MHD environment, large density fluctuations in both ions and neutrals lead to variations in the Alfvén speed and ionization fraction. These variations cause the parallel and perpendicular neutral–ion decoupling scales to span over an extended range. Consequently, the overall coupling status between ions and neutrals can be categorized into three regimes: (i) a strongly coupled regime, where scales larger than the maximum decoupling length engender single-fluid behavior; (ii) a intermediately coupled regime, corresponding to intermediate scales where the neutral and ions can be either locally coupled or locally decoupled, , depending on the local physical conditions; and (iii) a fully decoupled regime, at scales below the minimum decoupling scale, in which neutrals evolve an independent hydrodynamic cascade while ions follow a separate MHD cascade if turbulent energy is injected at these small scales.
2.2 Neutral-ion collisional damping of MHD turbulence
Strongly coupled regime: In the strongly coupled regime, ions and neutrals move in unison, effectively forming a single fluid that transports the MHD turbulence. Both species share and sustain the cascading turbulent energy without any noticeable separation in their dynamics.
Intermediate coupling: Transitioning to the intermediately coupled regime, neutrals begin to decouple locally from the ions. Although ions continue to experience frequent collisions with nearby neutrals—extending down to the ion–neutral decoupling scale—the local decoupling of neutrals leads to a significant modification of the ion dynamics. The MHD turbulence in ions is strongly damped by the frictional forces arising from these collisions. In this regime, the hydrodynamic turbulence in neutrals is also modified and damped due to interactions with ions (Hu et al., 2024).
Fully decoupled regime: In the fully decoupled regime, the behavior of the two components diverges markedly. Ions remain subject to frequent collisions with the surrounding neutrals, which results in the damping of their MHD turbulent cascade. Conversely, neutrals develop an independent hydrodynamic cascade. This decoupling leads to a clear separation in the dynamical evolution: while the ion turbulence remains damped by collisional friction, the neutrals evolve under a purely hydrodynamic regime.
The damping of MHD turbulence leads to a marked reduction in fluctuations of density, velocity, and especially the magnetic field. This diminished fluctuations in the magnetic field curtail the efficiency of pitch angle scattering for CRs, thereby altering their transport properties along the magnetic field lines.
Figure 1: An example of the scattering diffusion and mirror diffusion. A particle with a Larmor radius of 10 cells and an initial pitch angle cosine of 0.25 is ejected into the simulation A4 ($M_{s}=1.24$ and $M_{A}=1.21$). The particle trajectory is plotted in the 3D volume rendering cube of normalized magnetic field strength. The trajectory is colored to show the particle’s evolution time: time increases from blue to red. In the regimes dominated by a magnetic field mirror, the particle is oscillating back and forth but gradually diffusion out of the mirror.
2.3 CRs’ perpendicular superdiffusion induced by wandering magnetic field lines
CRs’ propagation along the magnetic field lines can be separated into two regimes based on the relationship between their mean free path $\langle\lambda_{\parallel}\rangle$ (parallel to the magnetic field).
(1) When $\langle\lambda_{\parallel}\rangle>L_{\rm inj}$: the scattering of CRs is negligible, and they move along magnetic field lines in a ballistic manner. The divergence of magnetic field lines primarily determines their diffusion perpendicular to the mean magnetic field. Specifically, the magnetic field lines separate superdiffusively in space Lazarian & Vishniac (1999):
$\langle z^{2}\rangle\sim\frac{x^{3}}{27L_{\rm inj}}M_{A}^{4},~{}~{}{M_{A}\leq 1},$ (3)
$\langle z^{2}\rangle\sim\frac{x^{3}}{27L_{\rm inj}}M_{A}^{3},~{}~{}{M_{A}>1},$
where $x$ is the distance traced along magnetic field lines that are initially infinitely close to each other [^2] A peculiar feature of this law is that for sufficiently large $x$, the line separation can be independent of the initial line separation as discussed in Eyink et al. (2011)., while $\sqrt{\langle z^{2}\rangle}$ is the averaged separation of magnetic field lines perpendicular direction gained as the result of this tracing. Note that the dependence on $M_{A}$ is different for sub-Alfvénic and super-Alfvénic turbulence.
The magnetic field wandering imprints its law on the CRs moving along magnetic field lines (Xu & Yan, 2013). Those exhibit superdiffusion in the direction perpendicular to the magnetic field (Yan & Lazarian, 2008; Lazarian & Yan, 2014). This superdiffusion in space is broadly related to the Richardson dispersion of magnetic field lines in time Eyink et al. (2013), but is not equivalent to it. The relation between the two can be exemplified by considering CR moving ballistically along the wandering magnetic field lines with constant velocity. After time $t$, the dispersion of CR positions is:
$\langle z^{2}\rangle\sim\frac{t^{3}}{27L_{\rm inj}}M_{A}^{4},~{}~{}{M_{A}\leq 1},$ (4)
$\langle z^{2}\rangle\sim\frac{t^{3}}{27L_{\rm inj}}M_{A}^{3},~{}~{}{M_{A}>1},$
The functional dependence of $\sqrt{\langle z^{2}\rangle}$ on $t$ formally coincides with the law of the Richardson dispersion of magnetic field lines. However, the process described above is not the advection of magnetic field lines by turbulence described by Richardson diffusion. Indeed, in our example above, the magnetic field lines are assumed to be static.
Note that Eq. 4 is valid for strong turbulence, specifically for scales smaller than $l_{\rm tr}=L_{\rm inj}M_{A}^{2}$ when $M_{A}\leq 1$ or $l_{A}=L_{\rm inj}M_{A}^{-3}$ when $M_{A}>1$.  [^3] In the weak turbulence regime (i.e., scales larger than $l_{\rm tr}$ or $l_{A}$), turbulence is wave-like rather than anisotropic-eddy-like. In this regime, the separation of magnetic field lines follows a random walk, resulting in normal diffusion of CRs: $\langle z^{2}\rangle\sim tL_{\rm inj}M_{A}^{4},~{}~{}{M_{A}\leq 1},$ (5) $\langle z^{2}\rangle\sim tL_{\rm inj}M_{A}^{-3},~{}~{}{M_{A}>1}.$ For a detailed derivation of these equations, see Yan & Lazarian (2008); Lazarian & Yan (2014); Hu et al. (2022).
Run $M_{s}$ $M_{A}$ $\frac{\gamma_{\rm d}\rho_{i}}{v_{\rm inj}/L_{\rm inj}}$ $k_{\rm dec,\parallel}\frac{L_{\rm box}}{2\pi}$ $k_{\rm dec,\bot}\frac{L_{\rm box}}{2\pi}$ $l_{\rm dec,\parallel}$ [pc] $l_{\rm dec,\bot}$ [pc] $l_{\rm dis}$ [pc] $L_{\rm box}$ [pc] Coupling status
A1 1.00 0.87 $5\times 10^{3}$ $10^{4}$ $7\times 10^{5}$ $10^{-6}$ $1.4\times 10^{-8}$ $10^{-4}$ $10^{-2}$ strong coupling
A2 0.88 0.78 $5\times 10^{2}$ $10^{3}$ $2\times 10^{4}$ $10^{-5}$ $5\times 10^{-7}$ $10^{-4}$ $10^{-2}$ coupling
A3 0.82 0.74 $5\times 10^{1}$ $10^{2}$ $7\times 10^{2}$ $10^{-4}$ $1.4\times 10^{-5}$ $10^{-4}$ $10^{-2}$ intermediate coupling
A4 1.24 1.21 $5\times 10^{0}$ $10^{1}$ $2\times 10^{1}$ $10^{-3}$ $5\times 10^{-4}$ $10^{-4}$ $10^{-2}$ decoupling
B1 9.49 0.88 $5\times 10^{3}$ $10^{4}$ $7\times 10^{5}$ $10^{-3}$ $1.4\times 10^{-5}$ $10^{-1}$ $10^{1}$ strong coupling
B2 9.55 0.89 $5\times 10^{2}$ $10^{3}$ $2\times 10^{4}$ $10^{-2}$ $5\times 10^{-4}$ $10^{-1}$ $10^{1}$ coupling
B3 9.60 0.91 $5\times 10^{1}$ $10^{2}$ $7\times 10^{2}$ $10^{-1}$ $1.4\times 10^{-2}$ $10^{-1}$ $10^{1}$ intermediate coupling
B4 11.05 1.09 $5\times 10^{0}$ $10^{1}$ $2\times 10^{1}$ $10^{0}$ $5\times 10^{-1}$ $10^{-1}$ $10^{1}$ decoupling
Table 1: The setup of two-fluid simulations at a resolution of $1200^{3}$ is summarized here. The sonic Mach number ($M_{s}$) and Alfvénic Mach number ($M_{A}$) represent the instantaneous RMS values at each snapshot. The plasma compressibility is given by $\beta=2(M_{A}/M_{s})^{2}$. The ratio of the neutral-ion collisional frequency to the eddy turnover rate is expressed as $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, which characterizes the coupling strength. The theoretically expected parallel and perpendicular components of the neutral-ion decoupling wavenumber are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively, while the corresponding neutral-ion decoupling scales are given by $l_{\rm dec,\parallel}$ and $l_{\rm dec,\bot}$. The numerical dissipation scale is represented by $l_{\rm dis}$.
(2) When $\langle\lambda_{\parallel}\rangle<L_{\rm inj}$: the scattering of CRs along magnetic field lines is non-negligible, and their propagation becomes diffusive rather than ballistic, as numerically tested in (Hu et al., 2022). Parallel diffusion can be influenced predominantly by either pitch-angle scattering diffusion or bouncing from magnetic mirrors.
Magnetic mirrors created by changes of magnetic field strength were assumed to trap CR with sufficiently large pitch angles, preventing them from further propagation unless their pitch angle decreases through resonant scattering Cesarsky & Kulsrud (1973). However, a recent study in Lazarian & Xu (2021) suggested that in MHD turbulence, the perpendicular superdiffusion of turbulent magnetic field lines prevents CRs from becoming trapped between magnetic mirrors, as the return of CR to the initial mirror is improbable. Instead, CRs repeatedly bounce among various magnetic mirrors while diffusing along the local magnetic field lines, leading to a new type of diffusion mechanism termed ”mirror diffusion” (Lazarian & Xu, 2021). According to Lazarian & Xu (2021), the transition between the scattering and mirror diffusion regimes occurs at a critical cosine of pitch angle $\mu_{c}$. For CRs with a pitch angle cosine $\mu$ greater than $\mu_{c}$, parallel diffusion is governed primarily by resonant pitch angle scattering, resulting in scattering diffusion; conversely, for $\mu$ less than $\mu_{c}$, mirror diffusion prevails. The mirror interactions of CRs are strong as they reverse the CRs’ motion. Thus, mirror diffusion is typically slower than that induced by resonance scattering Lazarian & Xu (2021). An example of the scattering diffusion and the mirror diffusion is given in Fig. 1.
A particle with a Larmor radius of 10 cells and an initial pitch angle cosine of 0.25 was injected into simulation A4 (with $M_{s}=1.24$ and $M_{A}=1.21$). Initially, the particle entered a region dominated by a magnetic field mirror, which caused it to oscillate back and forth. Over time, the particle gradually diffuses out of this mirror, accompanied by changes in its pitch angle. Eventually, the particle became dominated by scattering diffusion and entered another magnetic mirror region. Both scattering and mirror diffusion contribute to the reduction in the CRs’ diffusion perpendicular to the mean magnetic field. In this context, the perpendicular diffusion can be expressed as follows (Lazarian & Yan, 2014; Hu et al., 2022):
$\langle z^{2}\rangle\sim\frac{D_{\|}t^{3/2}}{27L_{\rm inj}}M_{A}^{4},~{}~{}{M_{A}\leq 1}$ (6)
$\langle z^{2}\rangle\sim\frac{D_{\|}t^{3/2}}{27L_{\rm inj}}M_{A}^{3},~{}~{}{M_{A}>1}$
where $D_{\parallel}$ is the parallel diffusion coefficient. The transport of CRs is thus highly sensitive to the value of their mean free path $\langle\lambda_{\parallel}\rangle$, which is influenced by magnetic field fluctuations, for instance via pitch angle scattering.
Strongly coupled regime: In the strongly coupled regime, ions and neutrals develop MHD turbulence due to sufficient collision and energy transfer. Suppose magnetic field fluctuations are significant in the strongly coupled regime such that $\langle\lambda_{\parallel}\rangle<L_{\rm inj}$, CRs experience strong scattering and propagate in a superdiffusive manner. Their perpendicular diffusion follows the scaling described by Eq. 5, with $\sqrt{\langle z^{2}\rangle}\propto t^{3/4}$.
Intermediate coupling and fully decoupled regimes: When transitioning into the intermediately coupled or fully decoupled regimes, the damping of MHD turbulence occurs and magnetic field fluctuations are reduced, leading to an increase in $\langle\lambda_{\parallel}\rangle$. As a result, CRs shift into the $\langle\lambda_{\parallel}\rangle>L_{\rm inj}$ regime. In this regime, CRs experience insignificant scattering, and their superdiffusion shifts to the scaling described by Eq. 4, with $\sqrt{\langle z^{2}\rangle}\propto t^{3/2}$, which is driven by the Richardson diffusion of magnetic field lines. The transition between these regimes highlights the critical role of ions and neutrals coupling in determining CR propagation.
Figure 2: Left three panels: $x-z$ slices of ion velocity ($v_{i}$; top) and neutral velocity ($v_{n}$; bottom) at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the kinetic energy spectrum. The theoretically expected parallel and perpendicular components of the neutral-ion decoupling wavenumber are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, for comparison with Kolmogorov and Burgers turbulence scaling, respectively.
Figure 3: Left three panels: $x-z$ slices of ion density ($\rho_{i}$; top) and neutral density ($\rho_{n}$; bottom) at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the density spectrum. The neutral-ion decoupling wavenumber’s theoretically expected parallel and perpendicular components are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, for comparison with Kolmogorov and Burgers turbulence scaling, respectively.
3 Numerical methods
3.1 Two-fluid simulations
The scale-free 3D two-fluid simulations analyzed in this work were performed with the AthenaK code (Stone et al., 2024). We model a two-fluid magneto-fluid system consisting of an ionized fluid (ions plus electrons) and a neutral fluid (Hu et al., 2024). The simulation domain is discretized on a 12003 uniformly staggered grid. Under the assumptions of an isothermal equation of state with constant sound speed $c_{s}$ and periodic boundary conditions, we solve the ideal two-fluid MHD equations:
$\partial\rho_{i}/\partial t+\nabla\cdot\left(\rho_{i}\mathbf{v}_{i}\right)=0,$ (7)
$\partial\rho_{n}/\partial t+\nabla\cdot(\rho_{n}\mathbf{v}_{n})=0,$
$\partial(\rho_{i}\mathbf{v}_{i})/\partial t+\nabla\cdot\left[\rho_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{T}+\left(c_{s}^{2}\rho_{i}+\frac{B^{2}}{8\pi}\right)\mathbf{I}-\frac{\mathbf{B}\mathbf{B}^{T}}{4\pi}\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=\gamma_{\rm d}\rho_{n}\rho_{i}(\mathbf{v}_{n}-\mathbf{v}_{i})+\mathbf{f}_{i},$
$\partial(\rho_{n}\mathbf{v}_{n})/\partial t+\nabla\cdot\left[\rho_{n}\mathbf{v}_{n}\mathbf{v}_{n}^{T}+c_{s}^{2}\rho_{n}\mathbf{I}\right]=\gamma_{\rm d}\rho_{n}\rho_{i}(\mathbf{v}_{i}-\mathbf{v}_{n})+\mathbf{f}_{n},$
$\partial\mathbf{B}/\partial t-\nabla\times(\mathbf{v}_{i}\times\mathbf{B})=0,$
$\nabla\cdot\mathbf{B}=0,$
where $\rho$ and $\mathbf{v}$ denote the mass density and velocity of the ionized (subscript ”$i$”) and neutral (subscript ”$n$”) fluids, respectively, and $\mathbf{B}$ is the magnetic field. The drag coefficient $\gamma_{\rm d}$ couples ions and neutrals through the collisional exchange of momentum, while the stochastic forcing terms $\mathbf{f}_{i}$ and $\mathbf{f}_{n}$ drive turbulence in both fluids with weights proportional to their densities.
[…]
Figure 4: Left three panels: $x-z$ slices of magnetic field strength at three different neutral-ion coupling statuses: strong coupling, intermediate coupling, and decoupling. Right panel: the magnetic field spectrum. To guide the eye, the dashed and dash-dotted black lines represent power-law slopes of -5/3 and -2, respectively.
Figure 5: The parallel mean free path $\lambda_{\parallel}$ in the unit of turbulence injection scale $L_{\rm inj}$ as a function of $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, which is the ratio of the neutral-ion collisional frequency to the eddy turnover rate.
4 Results
4.1 Damping of velocity, density, and magnetic field fluctuations
In this section, we focus on studying the spatial distributions of velocity, density, and magnetic field, as well as their spectra. Their probability distribution functions are given in the Appendix. A.
4.1.1 Velocity distribution and kinetic energy spectrum
Strongly coupled regime: Fig. 2 presents 2D slices of the velocity fields and turbulent kinetic energy spectra for ions and neutrals, taken perpendicular to the mean magnetic field at the center of the simulation box. In both transonic and supersonic conditions, the velocity structures of ions (upper panel) and neutrals (lower panel) are highly similar in the strongly coupled regime, closely resembling the single-fluid approximation. For the transsonic case, the ion and neutral velocity spectra approximately follow the Kolmogorov scaling, exhibiting a spectral slope of -5/3. In contrast, the supersonic case yields steeper spectra with slopes close to -2, indicative of Burgers scaling. These steeper spectra result from energy dissipation associated with shocks
Intermediate coupling: In the intermediate coupling regime, the 2D velocity slices and spectra of ions and neutrals remain similar to each other, but both spectra become steeper than the Kolmogorov and Burgers scalings. This steepening indicates significant damping of the turbulent cascade. In this regime, neutrals begin to decouple from ions, although the decoupling is incomplete due to the fluctuations in ionization fractions. Consequently, neutrals cannot fully establish an independent hydrodynamic cascade, resulting in persistently steep spectra  [^4] The reduced value of $\gamma_{\rm d}$ and a high ionization fraction of 0.1 may cause enhanced ion inertial and frictional damping. However, a low ionization fraction 0.01 test has been given in the Appendix in Hu et al. (2024), showing that the neutral kinetic energy spectrum remains steep in the intermediate coupling regime.. This behavior suggests that neutral-ion decoupling occurs gradually over a range of scales rather than sharply at a specific scale. Compared to the strongly coupled regime, the velocity distributions of both neutrals and ions exhibit a clear deficiency of small-scale structures.
In addition to the reason that fluctuations in ionization fraction result in different local decoupling scale, this regime is also likely to have a contribution from the anisotropy of MHD turbulence, which means Alfvén waves propagate predominantly along the magnetic field, while the turbulent cascade develops preferentially in the perpendicular direction (Goldreich & Sridhar, 1995; Lazarian & Vishniac, 1999). As a result, neutrals may decouple from ions along the parallel direction yet remain coupled in the perpendicular direction due to turbulent motion. This behavior implies that neutral-ion coupling has both parallel and perpendicular components. Consequently, the decoupling scale is anisotropic, with its parallel component potentially much larger than the perpendicular component (Xu et al., 2015; Hu et al., 2024). In contrast, the classical model of neutral-ion coupling (see Eq. 1) considers only Alfvén waves and thereby omits the perpendicular coupling.
Decoupled regime: In the decoupled regime, the ion and neutral spectra remain steep up to a wavenumber of approximately 30, which is close to the perpendicular decoupling scale. Beyond this scale, the neutral spectrum becomes shallower, indicating that neutrals have fully decoupled from ions and developed an independent hydrodynamic turbulent cascade characterized by a Kolmogorov slope. This trend was not observed in the earlier study by Hu et al. (2024) due to limited numerical resolution. Meanwhile, ions experience frequent collisions with neutrals, effectively damping ion turbulence. We find that the damping is more severe in the transonic case with the even steeper spectra slope than that in the supersonic case. It is likely that shock compression locally increases the density fluctuations and ionization fraction, efficiently reducing the decoupling. Although the velocity slices of ions and neutrals exhibit noticeable differences, both species predominantly retain only large-scale fluctuations.
4.1.2 Density distribution and density spectrum
Fig. 3 presents 2D density slices and corresponding density spectra for ions and neutrals. Under supersonic conditions, the density structures of ions (upper panel) and neutrals (lower panel) remain broadly similar across all coupling regimes, although subtle differences emerge. As decoupling progresses, small-scale density structures become less pronounced. In the strongly coupled regime, the neutral density spectra closely resemble those of ions, with both spectra exhibiting slopes shallower than the Kolmogorov scaling due to shock-induced small-scale density fluctuations. As the coupling weakens, the spectral slopes gradually steepen, particularly for neutrals, reflecting increased damping of small-scale density fluctuations.
In contrast, under transonic conditions, the density spectra exhibit distinct characteristics. In the strongly coupled regime, both ion and neutral density spectra generally follow the Kolmogorov scaling, similar to their velocity spectra. However, as decoupling occurs, the ion density structures diverge significantly from those of neutrals, despite the continued similarity in their velocity fields. Ion density distributions exhibit pronounced small-scale striations, which are absent in the neutral density distributions. Correspondingly, the ion density spectrum remains shallow, whereas the neutral density spectrum steepens noticeably. This difference likely arises because the velocity field is dominated by incompressible Alfvénic turbulence, while compressible turbulent motions primarily drive density fluctuations. Although neutrals remain coupled to Alfvénic turbulent motions, their coupling to compressible MHD turbulence weakens, preventing the formation of small-scale density structures induced by compressible motions. In the fully decoupled regime, filamentary structures in the ion density vanish, and the ion density spectrum steepens significantly beyond a wavenumber of approximately 10. The observed differences between supersonic and transonic conditions are most likely attributable to shocks, as shock compression enhances density fluctuations and increases the ionization fraction, thereby strengthening ion-neutral coupling.
Figure 6: Plots of perpendicular displacement of the particles $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}/L_{\rm inj}$ versus the CRs’ gyro periods $t\cdot\Omega$ in transonic $M_{s}\approx 1$ conditions. $l_{\rm diss}$ is the numerical dissipation scale and $L_{\rm inj}$ is the injection scale. The panels (a), (b), and (c) correspond to initial cosine of pitch angle $\mu\approx 1$, $\mu\approx 0.25$, and $\mu\approx 0.1$, respectively.
4.1.3 Magnetic field distribution and magnetic field spectrum
Fig. 4 presents 2D slices of the total magnetic field strength and corresponding magnetic energy spectra. In the strongly coupled regime, magnetic field fluctuations exhibit filamentary structures under transonic conditions, closely resembling the velocity structures shown in Fig. 2. The magnetic energy spectrum in this case generally follows a Kolmogorov scaling with a slope of approximately -5/3. Under supersonic conditions, numerous discontinuities associated with shocks are evident, resulting in a steeper magnetic energy spectrum characterized by a slope close to -2. Small-scale magnetic field structures become significantly less prominent in the intermediate coupled and decoupled regimes. Correspondingly, the magnetic energy spectra steepen beyond a slope of -2 for both supersonic and transonic conditions, indicating that magnetic energy becomes increasingly concentrated at larger scales. This behavior naturally arises due to substantial damping caused by neutral-ion collisions, effectively suppressing magnetic fluctuations.
4.2 Damping of magnetic field fluctuations increases CRs’ mean free path
Damping caused by neutral-ion collisions effectively suppresses magnetic fluctuations at smaller scales, implying reduced pitch-angle scattering of CRs and consequently a larger mean free path. To quantify the mean free path of CR particles, we adopt the methodology described in Xu & Yan (2013). Specifically, we inject 1000 test particles at random initial positions, assigning initial cosine pitch angles ($\mu$) around 1, 0.25, and 0.1, respectively. The particle Larmor radius is set to $r_{L}=v/\Omega$ and is set to 10 cells, where $\Omega$ is the frequency of a particle’s gyromotion. We trace the change of pitch angle until it achieves 90 degrees. The corresponding averaged distance traveled along the local magnetic field is taken as the parallel mean free path.
In Fig. 5, we present the mean free path $\langle\lambda_{\parallel}\rangle$ as functions of the dimensionless coupling strength $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, defined as the ratio of the neutral-ion collisional frequency to the eddy turnover rate. Under both supersonic and transonic conditions, an increase in $\langle\lambda_{\parallel}\rangle$ corresponds to weaker coupling, consistent with the suppression of small-scale magnetic fluctuations. When $\mu\approx 1$, the parallel motion of CRs along magnetic field lines is primarily governed by the resonant pitch-angle scattering and undergoes scattering diffusion. We trace the change of pitch angle until it achieves 90 degrees. The corresponding averaged distance traveled along the local magnetic field is taken as the parallel mean free path. However, at smaller pitch angles, mirror diffusion is dominant, and $\langle\lambda_{\parallel}\rangle$ is determined by the characteristic size of magnetic mirrors with which mirroring particles predominantly interact (Lazarian & Xu, 2021). As mirroring CRs repeatedly reverse their direction of motion along magnetic field lines, mirror diffusion strongly enhances CR confinement, resulting in a reduced mean free path at small $\mu$ (Zhang & Xu, 2023).
4.3 CRs’ perpendicular superdiffusion
To investigate perpendicular superdiffusion, we simultaneously inject 50 beams of test particles randomly distributed within the simulation cube, with each beam containing 20 particles. The initial spatial separation between particles within each beam is set to 10 grid cells. Consistent with the previous setup, all particles have a Larmor radius of $r_{L}=10$ cells and an initial $\mu\approx 1$, $\mu\approx 0.25$, or $\mu\approx 0.1$, respectively. We trace particle trajectories along local magnetic field lines at each time step and interpolate these trajectories to measure the perpendicular separation $\delta\widetilde{z}$ between particle trajectories at identical times. The root-mean-square (rms) value $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ is taken as the perpendicular displacement of the particles.
Fig. 6 presents the correlation of $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ and the CRs’ gyro periods $t\Omega$ in transonic condition with initial $\mu\approx 1$. In these conditions, the parallel mean free path $\langle\lambda_{\parallel}\rangle$ exceeds the turbulence injection scale $L_{\rm inj}$ across all three coupling regimes. After surpassing the numerical dissipation scale, $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ exhibits a power-law growth with a slope of approximately. The results in supersonic conditions are similar, as studied by Hu et al. (2022). The growth of $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ is slower in the decoupling regime is slower due to the damping of small-scale magnetic fluctuations, which significantly reduces particle scattering.
We repeat this analysis for particles with initial $\mu\approx 0.25$ and present the results in Fig. 6. In this scenario, mirror diffusion and scattering effects become important, reducing $\langle\lambda_{\parallel}\rangle$. In the decoupled regime, where $\langle\lambda_{\parallel}\rangle>L_{\rm inj}$, the perpendicular displacement $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ again follows a power-law scaling with a slope of approximately 3/2. However, in the intermediately and strongly coupled regimes, where damping of MHD turbulence is less significant, the power-law slope gradually deviates from 3/2, approaching a slope closer to 3/4. The power-law slope does not perfectly align with 3/4 because trajectories are averaged over 1000 particles. Since $\langle\lambda_{\parallel}\rangle$ is not significantly smaller than $L_{\rm inj}$, some particles with larger mean free paths still contribute to a slope closer to 3/2. Although the perpendicular displacement increases more slowly in these cases, the CR perpendicular motion remains superdiffusive. Additionally, $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ exhibits oscillations before entering the superdiffusive regime. These oscillations, absent in the $\mu\approx 1$ case, as shown in Fig. 6, could result from large magnetic mirrors, where CRs require more than one bounce for escaping (see Fig. 1). In the decoupled regime, small-scale magnetic fluctuations are damped, causing CRs to reflect predominantly from large-scale magnetic mirrors, thus producing more pronounced oscillations in $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$. The effects of interaction with mirrors of different sizes require further studies. Nevertheless, when the initial $\mu\approx 0.1$, the CRs’ $\langle\lambda_{\parallel}\rangle$ is smaller than $L_{\rm inj}$ in the three coupling/decoupling regimes. The power-law slope of $\sqrt{\langle\delta\widetilde{z}\rangle^{2}}$ gradually deviates from 3/2, but is closer to 3/4. The oscillation induced by mirrors is still apparent.
5 Conclusion
In this work, we have examined the effects of partially ionized, turbulent environments on CR transport using high-resolution, 3D two-fluid simulations. By explicitly accounting for ion-neutral interactions, we have shown that neutral-ion collisions significantly alter turbulence dynamics, leading to important consequences for CR diffusion. Our main findings can be summarized as follows:
  1. 1.
    Neutral-ion decoupling and turbulence damping: As neutral-ion coupling weakens, we observe substantial damping of velocity, density, and magnetic field fluctuations at small scales, resulting in steeper turbulence spectra than classical Kolmogorov or Burgers scalings. Under supersonic conditions, shocks partially counteract this damping, sustaining density fluctuations and ionization enhancements, thereby influencing the coupling strength and turbulence structure.
  2. 2.
    Impact on CR parallel propagation: The damping of magnetic field fluctuations reduces pitch-angle scattering, substantially increasing the CR mean free path along the magnetic field direction. Importantly, mirror diffusion becomes increasingly significant at larger pitch angles, decreasing the parallel transport and increasing CRs confinement.
  3. 3.
    Perpendicular CR superdiffusion: We identify two distinct superdiffusive regimes of CR transport perpendicular to the mean magnetic field. In the strongly coupled regime, frequent scattering and mirror diffusion limit the parallel mean free path, and CR perpendicular transport follows a superdiffusive scaling of $\propto t^{3/4}$. Conversely, in the intermediately coupled and fully decoupled regimes, the suppression of small-scale magnetic fluctuations reduces scattering, leading to transport dominated by magnetic field line wandering, characterized by a scaling of $\propto t^{3/2}$.
Y.H. thanks Anatoly Spitkovsky for the helpful discussion. Y.H. acknowledges the support for this work provided by NASA through the NASA Hubble Fellowship grant No. HST-HF2-51557.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. A.L. acknowledges the support of NSF grants AST 2307840. This work used SDSC Expanse CPU and NCSA Delta CPU through allocations PHY230032, PHY230033, PHY230091, PHY230105, PHY230178, and PHY240183, from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Appendix A Histograms of velocity, density, and magnetic field
Figure 7: Histograms of ion velocity (red) and neutral velocity (blue) in transonic (top) and supersonic (bottom) conditions.
We present the ion and neutral velocity probability distribution functions (PDFs) in Fig. 7. In the strong coupling regime, the PDFs for ions and neutrals are nearly identical. Differences become apparent in the intermediate coupling regime, with the largest divergence occurring in the decoupling region. Here, the ion velocity distribution shifts slightly toward higher values and broadens. This broadening is most likely caused by the unequal correlation time of the turbulence driving. Specifically, the initial correlation timescale of the driving force is set equal to the Alfvén speed crossing time in the strong coupling cases. When neutrals decouple from ions, they no longer support the development of Alfvén waves, and the effective Alfvén speed increases due to the reduced ion density. Consequently, the fixed correlation time in the simulation becomes too large for ions, resulting in a slower turbulence cascade and higher ion velocities (Hu et al., 2024). This effect is most pronounced in the supersonic decoupling case, where the sonic Mach numbers for neutrals and ions are approximately 7 and 11, respectively, compared to around 1.07 and 1.24 in the transonic decoupling case.
Figure 8: Same as Fig. 7, but for the logarithmic mass density $\log_{10}(\rho/\langle\rho\rangle)$.
Figure 9: Histograms of magnetic field strength in transonic (left) and supersonic (right) conditions.
Fig. 8 shows the PDFs of the logarithmic mass densities of ions and neutrals, normalized by their respective mean values. Under supersonic conditions, the PDFs of ions and neutrals are similar, regardless of the coupling status, because shocks effectively enhance the coupling in the density field (see Fig. 3). In transonic conditions with strong coupling, the ion PDFs closely resemble those of neutrals. However, in the intermediate coupling regime, the ion PDFs become more dispersed while the neutrals’ PDFs narrow, indicating more pronounced density fluctuations in ions. In the decoupling regimes, the PDFs of neutrals narrow further, signifying even less significant density fluctuations.
Fig. 9 presents the PDFs of the magnetic field strength. In both transonic and supersonic conditions, the PDFs in the strong coupling regime are broader than those in the decoupling regimes, indicating that magnetic field fluctuations are significantly damped when decoupling occurs.